# T-SCEND: Test-time Scalable MCTS-enhanced Diffusion Model

### v2

@@ 4.3 Inference of T-SCEND
1:  Input: EBM $E_{\theta}(\cdot)$, Diffusion Steps $T$, MCTS denoising start step $t_{s}$, Number of initial noise of Best-of-N random search $L$, Maximum MCTS branch count $K$, Maximum MCTS rollout step $N_{r}$, A set of initial diffusion state $\{\bm{x}_{T}^{(k)}\ |\ k=1,...,L\}$ sampled from Gaussian noise.
2:  // Random search for $T\rightarrow t_{s}$
@@
$\begin{split}\bm{x}_{t^{\prime}-1}^{(k)}&=\sqrt{\bar{\alpha}_{t^{\prime}-1}}\frac{\bm{x}_{t^{\prime}}-\sqrt{1-\bar{\alpha}_{t^{\prime}}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t^{\prime}},t^{\prime}\right)}{\sqrt{\bar{\alpha}_{t^{\prime}}}}\\ &\quad+\sqrt{1-\bar{\alpha}_{t^{\prime}-1}-\sigma_{t^{\prime}}^{2}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t^{\prime}},t^{\prime}\right)+\sigma_{t^{\prime}}\bm{\epsilon}^{(k)},\end{split}$ (9)
(3) Simulation: We randomly select a child node and perform a random simulation of the MDP until reaching a terminal state. For the simulation of MCTS denoising, we use DDIMs (Song et al., ) for fast sampling to obtain $\hat{\bm{x}}_{0}(\bm{x}_{t^{\prime}-1}^{(k^{*})})$ from the randomly chosen child node state ${\bm{x}}_{t^{\prime}-1}^{(k^{*})}$, and then use $E_{\theta}(\hat{\bm{x}}_{0}(\bm{x}_{t^{\prime}-1}^{(k^{*})}))$ as the reward for backpropagation.
To summarize, our proposed paradigm, T-SCEND, enhances performance-energy consistency and ease of sampling through LRNCL loss and KL regularization, respectively (Sec. 4.2), while fully unlocking the test-time scalability of diffusion models via hMCTS denoising (Sec. 4.3).
5 Experiments
Table 1: Success rate across different grid sizes of Maze and various number of given entries for naïve inference ($N=1$) for comparison of the generalization ability of models obtained by different training methods. Let $M$ denotes the grid size of Maze and $D$ denotes the number of given digits in Sudoku. Original denotes the original training method in Du et al. (2024). T-SCEND tr. (ours) denotes the training method of T-SCEND, and all subsequent descriptions will follow this definition. Bold font denotes the best model, and underline denotes the second best model. The same markings are used in the tables below.
Maze success rate Sudoku success rate
Methods $M=6$ $M=8$ $M=10$ $M=12$ $M=15$ $D=33$ $D=29$ $D=25$ $D=21$ $D=17$
T-SCEND tr. w/o LRNCL 1.0000 0.9922 0.7734 0.5625 0.2500 0.4219 0.1719 0.0469 0.0078 0.0000
T-SCEND tr. w/o KL 1.0000 0.9844 0.6953 0.4375 0.2812 0.4219 0.2578 0.0391 0.0078 0.0000
T-SCEND tr. (ours) 1.0000 1.0000 0.9922 0.9141 0.6562 0.1953 0.1016 0.0078 0.0000 0.0000
In the experiments, we aim to answer the following questions: (1) What factors contribute to the failure of the direct scaling-up diffusion model? (2) Does the training of T-SCEND unleash the test-time scalability of diffusion model? (3) Can the inference method of T-SCEND effectively achieve scaling up during test time? To answer these questions, we conduct extensive experiments on two challenging reasoning tasks: Maze and Sudoku. In Section  5.1, we find that the naïvely trained diffusion model falls short of scaling up during test time and identify the key reasons as insufficient performance-energy consistency and adversarial sampling. To address these challenges, the training methods of T-SCEND demonstrate a significant impact on unbinding the scalability of the diffusion model in Section 5.2. As shown in Fig 1, with the model trained with T-SCEND training methods, hMCTS denoising can better release the scalability enabling generalization to much more challenging tasks during inference (Sec. 5.3).
In Maze experiments, the datasets are generated by Ivanitskiy et al. (2023), and the diffusion model is mainly trained with Maze sizes of up to $6\times 6$ while tested on harder datasets with size significantly larger than $6\times 6$ (Fig. 1). For Sudoku experiments, the basic setting is adopted from Du et al. (2024) where the diffusion model is trained on SAT-Net dataset with 31 to 42 given entries (Wang et al., 2019) and tested on harder RRN dataset with 17 to 34 given entries (Palm et al., 2018) with fewer given entries. All models and inference methods are evaluated with solving success rate. For Maze, successful solving means finding a continuous path from starting location to target location without any breaking or overlapping with the wall; for Sudoku, successful solving means filling all the missing numbers that exactly match the ground-truth, both of which are very stringent metrics. More experimental details, hyperparameter analysis, and result visualizations can be found in Appendix B, Appendix C, and Appendix F, respectively.
Table 2: Success rate on Maze with grid size 15 and Sudoku harder dataset for comparison of the model’s ability to scale up under random search with different training methods. Here, $N_{r}=N,K=N,L=N$.
Maze success rate Sudoku success rate
Original, Random search 0.0625 0.0469 0.0547 0.0781 0.1016 0.1094 0.0859 0.1641 0.2188 0.2344 0.2422 0.2656 0.2812
T-SCEND tr. w/o LRNCL, Random search 0.2500 0.4297 0.5000 0.5391 0.6016 0.6094 0.1172 0.2656 0.3125 0.3438 0.3750 0.3906 0.4141
T-SCEND tr. w/o KL, Random search 0.2812 0.4688 0.4922 0.5547 0.5859 0.6250 0.1562 0.2422 0.2578 0.2656 0.2656 0.2891 0.3047
T-SCEND tr. (ours), Random search 0.2656 0.5859 0.6406 0.6875 0.6562 0.7031 0.0703 0.2500 0.3281 0.3438 0.3516 0.3594 0.3906
All the training methods are compared with the original training pipeline in Du et al. (2024). LRNCL and KL represent the loss terms ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$, respectively in Eq. 6. T-SCEND tr. w/o LRNCL, T-SCEND tr. w/o KL, and T-SCEND tr. (ours) represent the three training methods of T-SCEND, the last one being the full version. The MCTS denoising and hMCTS denoising are primarily compared with the best-of-N random search with the same computational budget.
5.1 Why originally trained model fail to scale up during test time?
Table 3: Performance-energy consistency of random search on Maze with grid size 15 and Sudoku harder dataset to test the effect of LRNCL loss. Here, $L=N$. Details of consistency calculation can be found in Appendix A.
Maze performance-energy consistency Sudoku Performance-energy consistency
Methods $N$=11 $N$=21 $N$=41 $N$=81 $N$=161 $N$=11 $N$=21 $N$=41 $N$=81 $N$=161
Original, Random search 0.3094 0.2916 0.3129 0.3143 0.3103 0.3955 0.3930 0.4010 0.4010 0.3986
T-SCEND tr. w/o KL, Random search 0.5375 0.4316 0.4902 0.5006 0.4998 0.4574 0.4528 0.4457 0.4459 0.4470
To investigate why diffusion models trained via the original method fail on more complex tasks, even with increased inference budgets yielding minimal improvement, we evaluate models trained on simpler datasets against harder tasks. As shown in Table 1, models trained on simpler tasks with original training methods exhibit significantly poorer performance on more complex tasks. For the Maze task, models trained on maze sizes of $4\times 4$, $5\times 5$, and $6\times 6$ demonstrate a rapid decline in success rate as the maze size increases. Specifically, the success rate decreases from 100% at $6\times 6$ to 6% at $15\times 15$. A similar trend is observed in the Sudoku task: models trained on datasets with 31 to 42 given entries fail to achieve high success rates when tested on datasets with 17 to 34 given entries, with the success rate dropping to below 5% when the number of entries is reduced to 25, and complete failure when only 21 entries are provided.
Building upon this, we conduct best-of-N random search to examine whether test-time scaling can enhance the performance of the originally trained model, as described in Section 4.3, and the results are presented in Table 2 and Table 3. In Table 2, it is evident that, for the originally trained model, increasing the inference budget $N$ in the Maze experiment only yields a modest improvement, with the success rate increasing by at most 5%, reaching a peak of 10%. In the Sudoku experiment, even with an increased budget of $N=320$, the success rate remains below 30%. The results in Table 3 further reveal that the energy model predictions of the originally trained model exhibit substantial misalignment with true performance metrics, with many good samples assigned higher energy values. In the Maze experiment, the consistency between the energy model and true performance is only 30% while the consistency in Sudoku task is around 40%. These results indicate that the originally trained model suffers from insufficient performance-energy consistency. Additionally, as seen in Table 10, even when ground truth is used to evaluate multiple candidates generated via random search with the originally trained model, the model’s performance remains unsatisfactory, with the Maze success rate not exceeding 20% and the Sudoku success rate below 30%. This result demonstrates that the originally trained model is affected by adversarial sampling. Even with an increased computation budget $N$, the generated samples remain suboptimal or correspond to incorrect global optima, hindering the acquisition of high-quality samples that align with the correct global optimum.
Figure 3: Scalability of different approaches on Maze and Sudoku.
Maze success rate Sudoku success rate
Original, Random search 0.0625 0.0469 0.0547 0.0781 0.1016 0.1094 0.0859 0.1641 0.2188 0.2344 0.2422 0.2656 0.2812
T-SCEND tr. (ours), MCTS denoising (ours) 0.2656 0.5859 0.6641 0.7344 0.6875 0.7031 0.0703 0.0859 0.0938 0.0859 0.1094 0.1250 0.1328
T-SCEND tr. (ours), Random search 0.2656 0.5859 0.6406 0.6875 0.6562 0.7031 0.0703 0.2500 0.3281 0.3438 0.3516 0.3594 0.3906
T-SCEND tr. (ours), hMCTS denoising (ours) 0.2656 0.6406 0.7266 0.7969 0.8203 0.8828 0.0703 0.2891 0.3203 0.3516 0.3672 0.4062 0.4297
5.2 The effect of T-SCEND training methods
To enhance the performance-energy consistency and reduce adversarial sampling, as detailed in 4.2, we augmented the originally trained method with additional LRNCL loss and KL loss. As shown in Table 3, incorporating LRNCL loss into the training process significantly improves the performance-energy consistency compared to the original model, with a greater than 20% improvement in the Maze task and a 5% improvement in the Sudoku task. In Table 2, for the Maze task, even without any additional computation (i.e., $N=1$), models trained with KL loss achieve a 20% increase in solving success rate compared to the original model. Models trained with LRNCL loss show an improvement of over 20%. Augmented with random search, as the budget $N$ increases, the improvement for all three training methods becomes more pronounced. Models trained with KL loss show a maximum success rate increase of approximately 50%, those with LRNCL loss exceed a 50% improvement, and the combination of both KL loss and LRNCL loss results in a maximum increase of about 60%. In the Sudoku task, using the naive inference method with $N=1$, both KL loss and LRNCL loss yield significant improvements. For Random Search, as $N$ increases, models trained with all three T-SCEND methods surpass the original model at $N=21$, with the gap widening further as the budget grows, showing a maximum increase of over 10%. Furthermore, Table 10 shows that the T-SCEND-trained energy-based diffusion model outperforms DDPM (Ho et al., 2020) under the same architecture. These results demonstrate that the training methods of T-SCEND can effectively improve performance-energy consistency and reduce adversarial sampling to unlock the test-time scalability of diffusion models.
5.3 Test-time Scalability of T-SCEND
Using the diffusion model trained with T-SCEND training methods, we evaluate various inference approaches on the Maze task with a grid size of $15\times 15$ and on a more challenging Sudoku dataset. As shown in Table 4, models trained with additional LRNCL loss and KL loss of T-SCEND demonstrate that, in the Maze experiment, the MCTS denoising method slightly outperforms random search, while our hMCTS denoising yields a significantly higher success rate, with a maximum improvement of approximately 18% than random search. Moreover, as the budget $N$ increases, the performance gap between hMCTS denoising and random search widens. In the Sudoku experiment, hMCTS denoising also consistently outperforms random search, with a maximum improvement of 5%. As illustrated in the scaling curve in Fig. 3, hMCTS denoising with the model trained with additional LRNCL loss and KL loss shows a marked improvement in solving success rate compared to both other inference methods and hMCTS denoising with the originally trained model. At the same time, the rate of improvement in solving success rate for other inference methods clearly slows down compared to hMCTS denoising. These results provide strong evidence that our inference method can effectively scale up during test time, offering a clear advantage over random search.
6 Conclusion
In this work, we have introduced Test-time Scalable MCTS-enhanced Diffusion Model (T-SCEND), a novel framework that enhances the scalability of diffusion models in testing to tackle more complex tasks than in training. We demonstrate that naïve scaling of existing diffusion models yields marginal performance gains. To address this, we introduce two auxiliary training loss, ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$, to improve performance-energy consistency and reduce adversarial sampling while integrating hybrid Monte Carlo Tree Search (hMCTS) denoising to fully exploit the model’s scalability in inference. We have conducted extensive experiments on Maze and Sudoku to validate the efficacy of our method and believe that T-SCEND offers a robust solution for test-time scaling, unlocking the notable potential of diffusion models. We discuss limitations and future work in Appendix E.
Appendix A Related algorithms and metric caculation
A.1 Performance-energy Consistency
In this paper, performance-energy consistency refers to the consistency between the results evaluated using an energy model and those evaluated using real-world metrics for the same sample. Specifically, the consistency requires that good samples are assigned low energy, while poor samples are assigned high energy. Performance-energy consistency measures the proportion of element pairs that maintain the same relative order in both permutations $X$ and $Y$, where $X$ and $Y$ represent the index arrays obtained by sorting the original energy values $\mathbf{E}=(E_{1},E_{2},\dots,E_{N})$ and performance metric values $\mathbf{P}=(P_{1},P_{2},\dots,P_{N})$, respectively, in ascending order. In this paper, the energy values are calculated by energy model $E_{\theta}(x_{0})$ for samples ${\bm{x}}_{0}$. The performance metric values are calculated as the L2 distance between the generated samples ${\bm{x}}_{0}$ and the ground truth under the given condition.
where:
  • $\binom{N}{2}=\frac{N(N-1)}{2}$ is the total number of pairs $(i,j)$ with $i<j$,
A.2 Adversarial sampling
During the sampling process, energy optimization often gets trapped in local minima or incorrect global minima, making it difficult to escape and hindering the sampling of high-quality samples.
A.3 Negative Sample Generation
A.4 Linear-regression algorithm
Given three points $(x_{1},y_{1})$, $(x_{2},y_{2})$, and $(x_{3},y_{3})$, we wish to fit a line of the form  (Lane et al., 2003):
$y=kx+b$
The mean of the $x$-coordinates and the mean of the $y$-coordinates are:
$\bar{x}=\frac{1}{3}(x_{1}+x_{2}+x_{3}),\quad\bar{y}=\frac{1}{3}(y_{1}+y_{2}+y_{3})$
The slope $k$ of the best-fit line is given by the formula:
$b=\bar{y}-k\bar{x}$
The equation of the best-fit line is:
$\hat{y}=kx+b$
@@ Appendix B Details of experiments
B.1 Detais of Sudoku experiments
For Sudoku experiment, the dataset, model architecture, and training configurations are adopted from Du et al. (2024). We mainly use solving success rate to evaluate different models. Model backbone and training configurations can be found in Fig. 4 and Table 5, respectively. All the exploration hyperparameters $c$ are set as 100 for Sudoku task.
@@ B.1 Detais of Sudoku experiments
Figure 4: The model architecture for T-SCEND on Sudoku task. The energy value is computed using the L2 norm of the final predicted output similar to Du et al. (2023), while the output is directly used as noise prediction for the diffusion baseline.
@@
B.2 Details of Maze experiments
The details of maze experiments and model backbone are provided in Table 6 and Fig. 5, respectively. The key metric, the maze-solving success rate is defined as the proportion of model-generated paths that have no breakpoints, do not overlap with walls, and begin and end at the start and target points, respectively. Maze datasets are generated by Ivanitskiy et al. (2023), and detailed hyperparameter configurations are in Table 6. All the exploration hyperparameters $c$ are set as 100 for Maze task.
@@ B.2 Details of Maze experiments
Figure 5: The model architecture for T-SCEND on Maze task. The energy value is computed using the L2 norm of the final predicted output similar to Du et al. (2023), while the output is directly used as noise prediction for the diffusion baseline.
@@ Appendix C Performance sensitivity to hyperparameters
In this subsection, we analyze the impact of several hyperparameters on the experimental results. As shown in Table 8, the influence of different noise scales on the performance of various methods is presented. The hMCTS denoising and random search require a relatively larger noise scale to better expand the search space and improve final performance, while the diffusion model with naive inference performs best with a smaller noise scale. As demonstrated in Table 7 and Fig. 6, the effect of varying inner-loop optimization steps on the results is also analyzed. It can be observed that performance improves gradually with an increasing number of steps, and after 5 steps, the performance stabilizes and the improvement slows down. Therefore, we chose 5 inner-loop optimization steps for the Maze experiments in this paper.
@@
T-SCEND tr. (ours), Naive inference 0.0000 0.1562 0.2109 0.2734 0.2812 0.2734 0.2812 0.2969 0.2969 0.2969
@@
T-SCEND tr. (ours), hMCTS denoising (energy) 0.3828 0.4375 0.5312 0.6094 0.6562 0.6953 0.7031 0.7344 0.7734 0.7969
T-SCEND tr. (ours), naive inference 0.3125 0.2656 0.2578 0.2344 0.2422 0.2656 0.2578 0.2422 0.2500 0.2500
T-SCEND tr. (ours), Random search(energy) 0.3906 0.4453 0.5312 0.5703 0.5938 0.6328 0.6641 0.6719 0.6797 0.6562
@@ Appendix D Additional results
The parameter $t_{s}$ controls the proportion of the total inference budget allocated to MCTS denoising. When $t_{s}=9$, it means only MCTS denoising is used, while $t_{s}=0$ means only best-of-N random search is employed. For $90<t_{s}<$ 9, hMCTS denoising is applied. As shown in Table 9 and Fig. 7, there is a noticeable peak in model performance as $t_{s}$ varies.
@@
T-SCEND tr. (ours), hMCTS denoising (energy) 0.6562 0.6094 0.6641 0.7969 0.7969 0.6406 0.4922 0.4922 0.4609 0.4453
Table 10: Success rate of Random search for different training methods on Maze with grid size 15 and Sudoku harder dataset guided with ground truth accuracy. Untrained, Random search (gt) represents use ground truth to guide the random search. Here, $L=N$. Bold font denotes the best model.
@@
Untrained, Random search (gt) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Original, Random search (gt) 0.0625 0.1250 0.1094 0.1328 0.1719 0.1719 0.0859 0.1641 0.2188 0.2344 0.2422 0.2656 0.2969
DDPM, Random search (gt) 0.0312 0.1094 0.1587 0.1746 0.2031 0.2422 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0156
T-SCEND tr. w/o LRNCL, Random search (gt) 0.2500 0.5078 0.5938 0.6562 0.7109 0.7422 0.1094 0.2578 0.2969 0.3438 0.3750 0.3828 0.4219
Table 11: Success rate and element-wise accuracy of Random search for different training methods on Sudoku harder dataset guided with ground truth accuracy. Here, $L=N$. Bold font denotes the best model.
@@
DDPM, Random search, GT accuracy guided 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0156 0.5071 0.6089 0.6316 0.6492 0.6691 0.6881 0.6999
Original, Random search, GT accuracy guided 0.0781 0.1641 0.2188 0.2344 0.2422 0.2656 0.2812 0.6650 0.7731 0.7952 0.8036 0.8217 0.8347 0.8491
T-SCEND tr. w/o LRNCL, Random search, GT accuracy guided 0.1094 0.2578 0.2969 0.3438 0.3750 0.3828 0.4219 0.6442 0.7855 0.8096 0.8317 0.8466 0.8628 0.8854
Appendix E Limitations and future work
Our inference framework primarily relies on MCTS, which presents two key limitations: (1) limited compatibility with parallel computing, and (2) challenges in effectively evaluating node quality during the early stages of denoising. Future work could explore integrating alternative search strategies, such as those proposed by Wu et al. (2024b). Additionally, to enhance performance-energy consistency, we introduce linear-regression negative contrastive learning, which enforces a linear relationship between energy and the distance to real samples. Further investigation is needed to assess the broader implications of this constraint and explore alternative regularization approaches. Lastly, while our current implementation utilizes Gaussian noise for branching, other diffusion-based branching mechanisms remain an open area for exploration.
Appendix F Visualization of results
F.1 Visualization of Maze experiments
This section presents visualizations of the training in Fig. 8, test Maze data in Fig. 9, and samples generated by different methods in Fig. 10. In the visuals, black pixels denote walls, green represents the starting point, red represents the goal point, blue marks the solved path, and white represents the feasible area. All visualizations are based on a few representative samples. The results from the training and test sets clearly show that the tasks in the test set are notably more challenging than those in the training set. Visual comparisons of samples generated by different methods reveal that the originally trained model, regardless of the inference strategy, performs consistently worse than T-SCEND.

### v3

@@ A.4 Inference algorithm
1:  Input: EBM $E_{\theta}(\cdot)$, Diffusion Steps $T$, MCTS denoising start step $t_{s}$, Number of initial noise for Best-of-N $L$, Maximum MCTS branch count $K$, Maximum MCTS rollout step $N_{r}$, A set of initial diffusion state $\{\bm{x}_{T}^{(k)}\ |\ k=1,...,L\}$ sampled from Gaussian noise.
2:  // BoN for $T\rightarrow t_{s}$
@@ Appendix B Details of experiments
B.1 Core Metric Definition
For Maze, successful solving means finding a continuous path from starting location to target location without any breaking or overlapping with the wall; for Sudoku, successful solving means filling all the missing numbers that exactly match the ground-truth, both of which are very stringent metrics.
B.2 Detais of Sudoku experiments
For Sudoku experiment, the dataset, model architecture, and training configurations are adopted from [5]. We mainly use solving success rate to evaluate different models. Model backbone and training configurations can be found in Fig. 5 and Table 6, respectively. All the exploration hyperparameters $c$ are set as 100 for Sudoku task.
@@ B.2 Detais of Sudoku experiments
Figure 5: The model architecture for VFScale on Sudoku task. The energy value is computed using the L2 norm of the final predicted output similar to [18], while the output is directly used as noise prediction for the diffusion baseline.
@@
B.3 Details of Maze experiments
The details of maze experiments and model backbone are provided in Table 7 and Fig. 6, respectively. The key metric, the maze-solving success rate is defined as the proportion of model-generated paths that have no breakpoints, do not overlap with walls, and begin and end at the start and target points, respectively. Maze datasets are generated by [24], and detailed hyperparameter configurations are in Table 7. All the exploration hyperparameters $c$ are set as 100 for Maze task.
@@ B.3 Details of Maze experiments
Figure 6: The model architecture for VFScale on Maze task. The energy value is computed using the L2 norm of the final predicted output similar to [18], while the output is directly used as noise prediction for the diffusion baseline.
@@
B.4 Compute Resources
For training efficiency, we report the training time of our models along with the machine information (1 GPU with 80 GB of VRAM, CPU: 16). On the Maze (100,000 steps) environment, naive training took 2 hours, while incorporating LRNCL increased the training time to 4 hours. Adding KL regularization resulted in a 3-hour training time, and combining LRNCL and KL required 6 hours. For the Sudoku (200,000 steps) environment, naive training took 4 hours. Training with +LRNCL took 5 hours, +KL took 6 hours, and +LRNCL+KL required 7 hours.
@@ Appendix C Performance sensitivity to hyperparameters
In this subsection, we analyze the impact of several hyperparameters on the experimental results. As shown in Table 9, the influence of different noise scales on the performance of various methods is presented. The hMCTS denoising and BoN require a relatively larger noise scale to better expand the search space and improve final performance, while the diffusion model with naive inference performs best with a smaller noise scale. As demonstrated in Table 8 and Fig. 7, the effect of varying inner-loop optimization steps on the results is also analyzed. It can be observed that performance improves gradually with an increasing number of steps, and after 5 steps, the performance stabilizes and the improvement slows down. Therefore, we chose 5 inner-loop optimization steps for the Maze experiments in this paper.
@@
VFScale tr. (ours), Naive inference 0.0000 0.1562 0.2109 0.2734 0.2812 0.2734 0.2812 0.2969 0.2969 0.2969
@@
VFScale tr. (ours), hMCTS denoising (energy) 0.3828 0.4375 0.5312 0.6094 0.6562 0.6953 0.7031 0.7344 0.7734 0.7969
VFScale tr. (ours), naive inference 0.3125 0.2656 0.2578 0.2344 0.2422 0.2656 0.2578 0.2422 0.2500 0.2500
VFScale tr. (ours), BoN(energy) 0.3906 0.4453 0.5312 0.5703 0.5938 0.6328 0.6641 0.6719 0.6797 0.6562
@@ Appendix D Additional results
The parameter $t_{s}$ controls the proportion of the total inference budget allocated to MCTS denoising. When $t_{s}=9$, it means only MCTS denoising is used, while $t_{s}=0$ means only BoN is employed. For $90<t_{s}<$ 9, hMCTS denoising is applied. As shown in Table 10 and Fig. 8, there is a noticeable peak in model performance as $t_{s}$ varies.
@@
VFScale tr. (ours), hMCTS denoising (energy) 0.6562 0.6094 0.6641 0.7969 0.7969 0.6406 0.4922 0.4922 0.4609 0.4453
Table 11: Success rate of BoN for different training methods on Maze with grid size 15 and Sudoku harder dataset guided with ground truth accuracy. Untrained, BoN (gt) represents use ground truth to guide the BoN. Here, $L=N$. Bold font denotes the best model.
@@
Untrained, BoN (gt) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Original, BoN (gt) 0.0625 0.1250 0.1094 0.1328 0.1719 0.1719 0.0859 0.1641 0.2188 0.2344 0.2422 0.2656 0.2969
DDPM, BoN (gt) 0.0312 0.1094 0.1587 0.1746 0.2031 0.2422 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0156
VFScale tr. w/o LRNCL, BoN (gt) 0.2500 0.5078 0.5938 0.6562 0.7109 0.7422 0.1094 0.2578 0.2969 0.3438 0.3750 0.3828 0.4219
Table 12: Success rate and element-wise accuracy of BoN for different training methods on Sudoku harder dataset guided with ground truth accuracy. Here, $L=N$. Bold font denotes the best model.
@@
DDPM, BoN, GT accuracy guided 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0156 0.5071 0.6089 0.6316 0.6492 0.6691 0.6881 0.6999
Original, BoN, GT accuracy guided 0.0781 0.1641 0.2188 0.2344 0.2422 0.2656 0.2812 0.6650 0.7731 0.7952 0.8036 0.8217 0.8347 0.8491
VFScale tr. w/o LRNCL, BoN, GT accuracy guided 0.1094 0.2578 0.2969 0.3438 0.3750 0.3828 0.4219 0.6442 0.7855 0.8096 0.8317 0.8466 0.8628 0.8854
Appendix E Extension of hMCTS to Conventional Diffusion Models
The inference method hMCTS of VFScale is not limited to energy-based diffusion models. With an external verifier, it can be applied to conventional models that predict noise directly. We validated this on the Maze task by replacing the learned energy with MSE as the reward, confirming hMCTS’s compatibility. Its gains over best-of-N search from the Table 13 below further demonstrate broad applicability across diffusion models.
Table 13: Success rate on Maze (grid size 15) using a conventional noise-predicting diffusion model (DDPM) guided by an external ground-truth verifier. We compare the performance of hMCTS with BoN under different compute budgets ($N$). Here, the reward is defined as the negative MSE to the ground-truth solution. Results validate the compatibility of hMCTS with conventional diffusion and its improved efficiency over best-of-$N$ sampling.
Methods $N=1$ $N=11$ $N=21$ $N=41$ $N=81$ $N=161$
DDPM, hMCTS, Ground-truth guided 0.0312 0.1328 0.1406 0.1875 0.2188 0.2422
DDPM, BoN, Ground-truth guided 0.0312 0.1094 0.1587 0.1746 0.2031 0.2422
Appendix F Broader Impacts
This paper aims to advance machine learning—particularly diffusion-based generative models. While our improvements promise higher-quality AI-generated content, they also carry potential societal risks. Accordingly, we must remain vigilant to unintended negative outcomes and guard against any unethical or illicit applications of this technology.
Appendix G Limitations and future work
Our inference framework primarily relies on MCTS, which presents two key limitations: (1) limited compatibility with parallel computing, and (2) challenges in effectively evaluating node quality during the early stages of denoising. Future work could explore integrating alternative search strategies, such as those proposed by [28]. Additionally, to enhance performance-energy consistency, we introduce linear-regression negative contrastive learning, which enforces a linear relationship between energy and the distance to real samples. Further investigation is needed to assess the broader implications of this constraint and explore alternative regularization approaches. Lastly, while our current implementation utilizes Gaussian noise for branching, other diffusion-based branching mechanisms remain an open area for exploration.
Appendix H Visualization of results
H.1 Visualization of Maze experiments
This section presents visualizations of the training in Fig. 9, test Maze data in Fig. 10, and samples generated by different methods in Fig. 11. In the visuals, black pixels denote walls, green represents the starting point, red represents the goal point, blue marks the solved path, and white represents the feasible area. All visualizations are based on a few representative samples. The results from the training and test sets clearly show that the tasks in the test set are notably more challenging than those in the training set. Visual comparisons of samples generated by different methods reveal that the originally trained model, regardless of the inference strategy, performs consistently worse than VFScale.